Find the value if c if O Σα + c) = 2 m=2
Answer #1

Sola Ž Litejon = 2 n=2 n E (1tc) Beo= Ž (ic) n=2 n=2 It is a glomettic series with to Common ratio ㅗ IHC 2 first tom a (1+c)-> for C=-1-3 2 L-1 - convergent Geometric is not tia velue of a for or this Geometric series convergent for Irkl where Hence